OpenStudy (abbles):

What are the possible rational zeros of f(x) = 3x4 + x3 - 13x2 - 2x + 9?

OpenStudy (kc_kennylau):

Quartic equations. What. The. Hell.

OpenStudy (abbles):

I am confused on this one. According to Descartes' Rule of Signs, it has two rational zeros, correct? I got plus or minus 1 and plus or minus 3 but it says it's wrong.

OpenStudy (kc_kennylau):

you wanna try factoring it with the factor theorem?

OpenStudy (abbles):

sure?

OpenStudy (kc_kennylau):

Coz surely you can't factorize a quartic using formulas

OpenStudy (abbles):

quartic?

OpenStudy (kc_kennylau):

quartic as in linear quadratic cubic quartic quintic

OpenStudy (abbles):

okay well my lesson says to use the Rational Root Theorem

OpenStudy (abbles):

which states that p/q = possible positive rational zeros

OpenStudy (anonymous):

list all number of the form \(\frac{p}{q}\) where \(p\)divides 9, i.e. \(p=1,3,9\) and \(q\) divides \(3\) so \(q=1,3\) dont forget the \(\pm\)

OpenStudy (anonymous):

the list will look like \[\pm1,\pm3,\pm9,\pm\frac{1}{3},\pm\frac{1}{9}\] i think that is all of them

OpenStudy (abbles):

oh, can q also equal 1? I was under the impression that it had to equal 3.

OpenStudy (abbles):

@satellite73 do you think you can help me out with another problem real quick?